1) Balance the equation

a) __A2B3 + __C = __C3B4 + ___A

b) __D + __E2= __D2E3

c) __FGH2= __FG + __F2H

d) __I + __J2 = __IJ5


1
Expert's answer
2021-06-10T03:33:16-0400

a) "4A_2B_3 + 9C = 3C_3B_4 +8A"

b) "4D + 3E_2= 2D_2E_3"

c) "2FGH_2= 2FG + 2H_2"

d) "2I + 5J_2 = 2IJ_5"